OpenStudy (anonymous):

The Product of two Consecutive room numbers is 182. Find the room numbers. List number of first room and second.

OpenStudy (anonymous):

a * b = 182 b - a = 1 13, 14

OpenStudy (anonymous):

thank you, can you help me with 3 more? :D That would be really awesome. But you dont have to if you dont want to :)

OpenStudy (anonymous):

ill write it out as new question

myininaya (myininaya):

n and n+1 are consecutive numbers their product is 182 so that means we have n(n+1)=182 => \[n^2+n=182= > n^2+n-182=0\] \[(n+14)(n-13)=0\] \[=> n=-14 \text{ or } n=13\] But n=13 since 13 is the only positive number here and n+1=13+1=14 the page numbers are 13,14